# Exploration vs Exploitation: Rethinking RLVR through Clipping, Entropy, and Spurious Reward

### v2

@@ Appendix A Related Works
Direct alignment methods (Rafailov-2023-Direct) provide a simple and stable offline alternative to RLHF. Extensions to DPO include broader ranking objectives (Dong-2023-RAFT; Yuan-2023-RRHF; Song-2024-Preference; Chen-2024-Noise; Liu-2025-LiPO) and lightweight variants that remove the reference model (Hong-2024-ORPO; Meng-2024-SimPO). Since DPO avoids reward model, limited human preference data becomes a key bottleneck; recent work addresses this by generating additional preference pairs via SFT policies (Zhao-2023-Slic; Liu-2024-Statistical). The framework has also been generalized to token-level MDPs (Rafailov-2024-From) and broader RL settings (Azar-2024-General). Complementary approaches incorporate online human feedback to reduce distribution shift and improve reasoning (Dong-2024-RLHF; Xiong-2024-Iterative; Pang-2024-Iterative). Another line studies unintentional alignment and proposes remedies (Pal-2024-Smaug; Tajwar-2024-Preference; Liu-2024-Provably; Xiao-2024-Bias; Yuan-2025-Advancing; Razin-2025-Unintentional; Chen-2025-Compo). For example, Razin-2025-Unintentional filter noisy preference pairs using CHES similarity, while Chen-2025-Compo show that combining comparison oracles with DPO mitigates unintended alignment effects.
@@ Appendix B Additional Experimental Results
We further analyze the effect of varying the clipping ratio threshold $\varepsilon$. Indeed, Figure 1 examined the cases $\varepsilon=0.2$ and $\varepsilon=\infty$ (no clipping), showing that relaxing the clipping threshold does not degrade performance. However, this does not yet confirm robustness under stricter clipping. To address this, we report additional results for $\varepsilon\in{0.15,0.1}$ in Figure 5. Across these settings, we observe behavior consistent with Figure 1: (i) some runs fail to improve, as predicted by our probabilistic reward-misalignment framework; and (ii) successful runs converge to roughly 70% validation accuracy regardless of the clipping strength. As discussed in §4.1, clipping primarily influences policy entropy. Among the improving trials, stricter clipping tends to reduce variance across seeds, reflecting a more deterministic policy toward convergence. Taken together, these results indicate that our findings remain robust under different choices of $\varepsilon$.
@@
We empirically evaluate the skewness measure $\Phi(\pi)$ introduced in Remark 4.2 on the actual Qwen-Math-7B policy. Recall that under unclipped training, entropy can increase after a single update only when $\Phi(\pi)<0$. Since the policy induces different action distributions $\pi(a\mid\mathbf{x})$ for different input questions $\mathbf{x}$, we estimate skewness across questions by sampling the first 500 examples from the DeepScaleR training set (Luo-2025-Deepscaler). For each question $\mathbf{x}$, we generate 64 responses $\mathbf{y}$ from the policy using the same sampling and decoding hyperparameters as in §3.2 and compute an empirical estimate of $\Phi(\pi(\cdot\mid\mathbf{x}))$. We visualize selected prompts $\mathbf{x}_{i}$ along with their corresponding skewness values in Figure 8, providing a clearer picture of how Qwen-Math-7B behaves across the dataset. Among 500 sampled questions, 358 ones satisfy $\Phi(\pi(\cdot\mid\mathbf{x}_{i}))<0$, which is consistent with the observed entropy increases for the unclipped training.
@@ Appendix C Theoretical Analysis
Following Williams-1992-Simple and Li-2024-Remax, we define the clipped surrogate loss with per-token ratios without adding separate length normalization terms as follow,
@@
and evaluate at the realized pairs $(a,\mathbf{h})=(\mathbf{y}^{(i)}_{t},\mathbf{h}^{(i)}_{t})$ in training. Note that the above GRPO update can be analyzed by interpreting it as one natural policy gradient (NPG) step under softmax tabular parametrization (Agarwal-2021-Theory).
@@
Interpreting the GRPO update as a natural policy gradient (NPG) step has been widely adopted to study entropy dynamics throughout training (Cui-2025-Entropy). Here, we summarize the key components of GRPO in Algorithm˜1, which motivate our reduction to an NPG-style update for analyzing the effect of clipping in GRPO. We note that Algorithm˜1 should be viewed as an abstraction of GRPO implementations used in practice (Shao-2024-Deepseekmath).
In the outer loop, a reference policy is fixed once per iteration, and the per-step objective may include a KL penalty that constrains the updated policy $\pi_{\theta}$ to remain close to $\pi_{\textnormal{ref}}$, thereby controlling the effective step size and preventing excessive policy drift. Recent “zero-RL" setups (see Yu-2025-DAPO), which are also adopted in the empirical evaluation of Shao-2025-Spurious, set the KL coefficient to zero, effectively removing the explicit KL term from the objective. As such, we likewise omit the KL term in our analysis. Under this regime, the outer loop does not affect the subsequent analysis.

### v3

@@ Appendix A Related Works
Direct alignment methods (Rafailov et al., 2023) provide a simple and stable offline alternative to RLHF. Extensions to DPO include broader ranking objectives (Dong et al., 2023; Yuan et al., 2023; Song et al., 2024; Chen et al., 2024a; Liu et al., 2025a) and lightweight variants that remove the reference model (Hong et al., 2024; Meng et al., 2024). Since DPO avoids reward model, limited human preference data becomes a key bottleneck; recent work addresses this by generating additional preference pairs via SFT policies (Zhao et al., 2023; Liu et al., 2024a). The framework has also been generalized to token-level MDPs (Rafailov et al., 2024) and broader RL settings (Azar et al., 2024). Complementary approaches incorporate online human feedback to reduce distribution shift and improve reasoning (Dong et al., 2024; Xiong et al., 2024; Pang et al., 2024). Another line studies unintentional alignment and proposes remedies (Pal et al., 2024; Tajwar et al., 2024; Liu et al., 2024b; Xiao et al., 2024; Yuan et al., 2025; Razin et al., 2025; Chen et al., 2025b). For example, Razin et al. (2025) filter noisy preference pairs using CHES similarity, while Chen et al. (2025b) show that combining comparison oracles with DPO mitigates unintended alignment effects.
@@ Appendix B Additional Experimental Results
We further analyze the effect of varying the clipping ratio threshold $\varepsilon$. Indeed, Figure 1 examined the cases $\varepsilon=0.2$ and $\varepsilon=\infty$ (no clipping), showing that relaxing the clipping threshold does not degrade performance. However, this does not yet confirm robustness under stricter clipping. To address this, we report additional results for $\varepsilon\in\{0.15,0.1\}$ in Figure 5. Across these settings, we observe behavior consistent with Figure 1: (i) some runs fail to improve, as predicted by our probabilistic reward-misalignment framework; and (ii) successful runs converge to roughly 70% validation accuracy regardless of the clipping strength. As discussed in §4.1, clipping primarily influences policy entropy. Among the improving trials, stricter clipping tends to reduce variance across seeds, reflecting a more deterministic policy toward convergence. Taken together, these results indicate that our findings remain robust under different choices of $\varepsilon$.
@@
We empirically evaluate the skewness measure $\Phi(\pi)$ introduced in Remark 4.2 on the actual Qwen-Math-7B policy. Recall that under unclipped training, entropy can increase after a single update only when $\Phi(\pi)<0$. Since the policy induces different action distributions $\pi(a\mid\mathbf{x})$ for different input questions $\mathbf{x}$, we estimate skewness across questions by sampling the first 500 examples from the DeepScaleR training set (Luo et al., 2025). For each question $\mathbf{x}$, we generate 64 responses $\mathbf{y}$ from the policy using the same sampling and decoding hyperparameters as in §3.2 and compute an empirical estimate of $\Phi(\pi(\cdot\mid\mathbf{x}))$. We visualize selected prompts $\mathbf{x}_{i}$ along with their corresponding skewness values in Figure 8, providing a clearer picture of how Qwen-Math-7B behaves across the dataset. Among 500 sampled questions, 358 ones satisfy $\Phi(\pi(\cdot\mid\mathbf{x}_{i}))<0$, which is consistent with the observed entropy increases for the unclipped training.
@@ Appendix C Theoretical Analysis
Following Williams (1992) and Li et al. (2024b), we define the clipped surrogate loss with per-token ratios without adding separate length normalization terms as follow,
@@
and evaluate at the realized pairs $(a,\mathbf{h})=(\mathbf{y}^{(i)}_{t},\mathbf{h}^{(i)}_{t})$ in training. Note that the above GRPO update can be analyzed by interpreting it as one natural policy gradient (NPG) step under softmax tabular parametrization (Agarwal et al., 2021).
@@
Interpreting the GRPO update as a natural policy gradient (NPG) step has been widely adopted to study entropy dynamics throughout training (Cui et al., 2025). Here, we summarize the key components of GRPO in Algorithm˜1, which motivate our reduction to an NPG-style update for analyzing the effect of clipping in GRPO. We note that Algorithm˜1 should be viewed as an abstraction of GRPO implementations used in practice (Shao et al., 2024).
In the outer loop, a reference policy is fixed once per iteration, and the per-step objective may include a KL penalty that constrains the updated policy $\pi_{\theta}$ to remain close to $\pi_{\textnormal{ref}}$, thereby controlling the effective step size and preventing excessive policy drift. Recent “zero-RL" setups (see Yu et al., 2025), which are also adopted in the empirical evaluation of Shao et al. (2025), set the KL coefficient to zero, effectively removing the explicit KL term from the objective. As such, we likewise omit the KL term in our analysis. Under this regime, the outer loop does not affect the subsequent analysis.